OpenStudy (anonymous):

^3 sqr rt 512x^5/y^3

OpenStudy (anonymous):

put that in the equation section below. what u said doesnt make sense

OpenStudy (anonymous):

\[\frac{\sqrt[3]{512x^5}}{y^3}????\]

OpenStudy (anonymous):

ohh that makes sense

OpenStudy (anonymous):

can i answer mathematical?? this is the first one i know! haha

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

ok the answer is (hold on let me put it on equation)

OpenStudy (anonymous):

\[\sqrt[8x]{x^2}/y\]

OpenStudy (anonymous):

\[\sqrt[3]{\frac{512x^5}{y^3}}=\sqrt[3]{\frac{8^3*x^3*x^2}{y^3}}=\frac{8x}{y}\sqrt[3]{x^2}\]